OpenStudy (anonymous):

What is the solution to the equation? t(9 – 1) = 5 x 9 – 5 x 1

OpenStudy (anonymous):

9t - t = 45 - 5 8t = 40. Therefore, t = 5.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome. =]